Given the following production plan: Briefly explain the differences between a Chase Production Strategy and a Level Production Strategy, then use: (a) chase production strategy and (b) level production strategy to compute the monthly production, ending inventory or backlog, and workforce levels. A worker is capable of producing 50 units per month. Assume the beginning inventory as of January is 500 units, and the firm desires to have 200 units inventory at the end of June.

Demand production EI workforce - Month- 2000 Jan- 3000 Feb Mar- 5000 Apr-6000 May-6000 Jun-2000
Answer #1

A chase production strategy involves producing the exact demand quantities by hiring and firing employees as required. In this strategy, the goal is to avoid inventory accumulation and keep the capacity exactly equal to the demand for all the periods. A level strategy, on the other hand, focuses on leveled resources and the same production output in every period. The protection against the fluctuation of demand is covered by the inventory built up or carried forward from the previous periods.

(a)

Chase plan:

Jan Feb Mar Apr May Jun
Demand 2,000 3,000 5,000 6,000 6,000 2,000
Production 1,500 3,000 5,000 6,000 6,000 2,200
Ending Inventory 500 0 0 0 0 0 200
Workforce 30 60 100 120 120 44

Calculations:

B C D E F G H Jan Feb Mar Apr May 2000 3000 5000 6000 6000 2000 1500 3000 5000 6000 2200 =C5+D4-D3 =D5+E4-E3 =E5+F4-F3 =F5+G4

(b)

Level plan:

Total demand = 24,000

Net production required = 24000 - beginning inventory + ending inventory = 24000 - 500 + 200 = 23,700

So, monthly production level = 23700 / 6 = 3,950

Jan Feb Mar Apr May Jun
Demand 2,000 3,000 5,000 6,000 6,000 2,000
Production 3,950 3,950 3,950 3,950 3,950 3,950
Ending Inventory 500 2,450 3,400 2,350 300 -1,750 200
Workforce 79 79 79 79 79 79

As we can see, a backorder of 1,750 units will occur in May which will be recovered in the next month.